OpenStudy (anonymous):

i have not done LCD for 7 years. im still reading about it and i still dont get it. can some one help me?

OpenStudy (anonymous):

The LCD of two numbers is the smallest number that both multiply into by some other number.

OpenStudy (anonymous):

ok. so what would the lcd of 20 be? for example

OpenStudy (anonymous):

Find an example in a book. It usually involves two numbers

OpenStudy (anonymous):

LCD is between two or more numbers.

OpenStudy (anonymous):

LCD of 5 and 3 is 15, because that is the smalled number both 5 and 3 go into.

OpenStudy (anonymous):

ok. i get it now. thanks so much:)

OpenStudy (anonymous):

here is how you find it for big numbers: factor each number into prime factors, then take each prime to the highest power you see in any one number. 5 and 3 are both primes, so just multiply. but for say 8 and 12 you would write \[8=2^3\] \[12=2^2\times 3\] so least common multiple is \[2^3\times 3 = 24\] of course you might have just seen the 24, but for 84 and 98 you would probably not see the answer. \[84=2^2\times 3 \times 7\] \[98=2 \times 7^2\] least common multiple is \[2^2 \times 3 \times 7^2 = 588\]